OpenStudy (anonymous):

$$\begin{align}\mathsf{\text{Compute}\;\lim_{n\to \infty} \left(a_n\right)^{-1/n^2} \text{where,}\, \\ a_n=\left(1+\frac{1}{n^2}\right)\cdot\left(1+\frac{2^2}{n^2}\right)^2\cdots\left(1+\frac{n^2}{n^2}\right)^{n}}.\end{align}$$

OpenStudy (anonymous):

Latex isn't processing? Sorry.

OpenStudy (anonymous):

Try refreshing.

OpenStudy (experimentx):

Jeez ... where do you get these questions??

OpenStudy (anonymous):

Hmm Q37. lol

OpenStudy (experimentx):

\[ \huge e^{\frac{-1}{n^2}(\ln((1+\frac{1}{n^2})(1+\frac{2^2}{n^2})^2....(1+\frac{n^2}{n^2})^n)}\]

OpenStudy (anonymous):

Yeah, that's what I did too.

OpenStudy (experimentx):

thanks ... i'll save a copy of it!!

OpenStudy (experimentx):

\[ \huge e^{\frac{-1}{n^2}(\ln(1+\frac{1}{n^2})+\ln(1+\frac{2^2}{n^2})^2+....\ln(1+\frac{n^2}{n^2})^n)}\]

OpenStudy (experimentx):

\[ \LARGE e^{\frac{-1}{n^2}(\ln(1+\frac{1}{n^2})+2\ln(1+\frac{2^2}{n^2})+....+n\ln(1+\frac{n^2}{n^2}))}\] Looks like same kinda problem we faced before .. there was factorial before ... now we have some other thing!!

OpenStudy (experimentx):

try expanding log inside and ... make it more ulgy http://www.wolframalpha.com/input/?i=expand+ln%281%2Bx%29+at+0 .... i'm looosing drive!!!

OpenStudy (anonymous):

\[\small-\frac{1}{n^2}\left(\ln\left(n^2+1^2\right) - \ln n^2 + 2\ln\left(n^2 + 2^2\right)-2\ln n^2 + \ldots +n\ln\left(n^2+n^2\right) - n\ln n^2 \right)\]

OpenStudy (anonymous):

But I don't think it's of much help.

OpenStudy (experimentx):

let's try wolframing!!

OpenStudy (anonymous):

lol

OpenStudy (experimentx):

i don't know even how to ask wolfram .. lol

OpenStudy (anonymous):

Can't we use the inequality here? \[\small \int_1^nx\ln \left(n^2+x^2\right) \le \sum_{x=1}^nx\ln\left(n^2+x^2\right)\le \int_1^nx\ln\left(n^2+x^2\right)+\ln\left(n^2+n^2\right)\]I am not sure, this maybe a fallacy.

OpenStudy (anonymous):

Neither do I lol :/

OpenStudy (anonymous):

$$\small \int_1^nx\ln \left(n^2+x^2\right) \le \sum_{x=1}^nx\ln\left(n^2+x^2\right)\le \int_1^nx\ln\left(n^2+x^2\right)+n\ln\left(n^2+n^2\right)$$

OpenStudy (experimentx):

you sure about this Riemann sum??

OpenStudy (anonymous):

I am not :/

OpenStudy (anonymous):

Hmm I will come back in an hour or so. I am not on laptop, netbook :/ and this site is lagging alot. Goodluck.

OpenStudy (experimentx):

haha ... don't expect anything from me!! if i'm lucky ... then consider yourself lucky!!

OpenStudy (experimentx):

After hell lotta simplification ..... \[ \huge e^{\frac{-1}{n^2}(\frac{n^2}{2} (\ln 4 - 1))} = e^{\frac12 - \ln2} \]